What is the value of a real estate property if the NOI next year is projected to be $100,000, the discount rate or opportunity cost of capital is 12%, and the NOI is expected to grow by 2% per year?
Answer #1

HI

This is the example of perpetuity stream.

Where NOI next year N1= $100,000

discount rate k = 12%

growth rate g = 2%

So value of real estate as per perpetuity formula = N1/(k-g)

= 100000/(0.12-0.02) = 100000/0.1

= $1,000,000

Thanks